# Complementarity in Human-AI Collaboration: Concept, Sources, and Evidence

### v1

@@ 5.5 Future Work
There are several potential areas of future research on human-AI complementarity. The conceptualization could obviously be applied to other domains in the future, but there are also methodological avenues to pursue.
Future work should expand our knowledge about relevant sources of complementarity potential. In this work, we have shown that information and capability asymmetry could be promising sources of complementarity potential. Regarding information asymmetry, we experimentally evaluated unique human contextual information. Investigating unique AI contextual information in future work could also produce interesting insights. It would, moreover, be worthwhile developing criteria regarding the utility of contextual information and the degree of capability asymmetry.
Furthermore, team settings could comprise more than two team members, considering multiple AI models or multiple humans. Team design principles could be derived on how to ensure a sufficient amount of complementarity potential and on how to select and combine human and artificial team members.
Different collaboration mechanisms could also be developed and evaluated. While we have focused on humans handling the integration in both experimental studies, the AI might conceivably also undertake the integration. Alternatively, decisions might not be integrated, but task instances might be delegated to a team member deciding on behalf of the team, with the delegation initiated by either the human or the AI (Fügener et al.,, 2022). Future work could investigate these collaboration forms in terms of the realization of complementarity potential and their suitability for specific use cases and application domains.
@@ 6 Conclusion
So far, human-AI decision-making has been primarily concerned with AI systems helping human users. However, since the number of decision tasks that can be automated (i.e., can be solved by the AI alone) is increasing steadily, the focus has shifted to the purposeful design of the collaboration between humans and AI as team members—thereby shaping the future of work with AI. The ultimate objective of these teams must be to achieve complementary team performance ($CTP$), with the team outperforming each individual team member. The IS community is predestined to drive the development of appropriate theories and to lay the foundation for practical applications. We hope that the conceptual foundation developed in this paper will provide fruitful ground for future research, and that the empirical studies illustrate the validity and potential of the human-AI complementarity paradigm.

### v2

@@ 5.5 Future Work
There are several potential areas for future research on human-AI complementarity. The conceptualization could obviously be applied to other domains in the future, but there are also methodological avenues to pursue.
Future work should expand our knowledge about relevant sources of complementarity. In this work, we have shown that information and capability asymmetry are the key sources of complementarity, but we have not specified them further. Regarding information asymmetry, we experimentally evaluated unique human contextual information. Investigating unique AI contextual information in future work could also produce interesting insights. Moreover, it would be worthwhile developing criteria regarding the utility of contextual information and the degree of capability asymmetry.
Furthermore, team settings could comprise more than two team members, considering multiple AI models or multiple humans. Team design principles could be derived on how to ensure a sufficient degree of complementarity and on how to select and combine human and artificial team members (Hemmer et al., 2022).
Different collaboration forms and mechanisms could also be developed and evaluated. While we have focused on humans handling the integration in both experimental studies, the AI might conceivably also undertake the integration. Alternatively, decisions might not be integrated, but task instances might be delegated to a team member deciding on behalf of the team, with the delegation initiated by either the human or the AI (Fügener et al., 2022). Future work could investigate these collaboration forms in terms of their suitability for specific use cases and application domains.
Finally, although human-AI decision-making is an important application of human-AI collaboration, other types of problems, e.g., creative or generative tasks (Schmidt et al., 2023), may also receive scrutiny with regard to complementarity (Malone et al., 2023).
@@ 6 Conclusion
So far, human-AI collaboration in decision-making has been primarily concerned with AI systems helping human users. However, since the number of decision tasks that can be automated (i.e., can be solved by the AI alone) is increasing steadily, the focus has shifted to the purposeful design of the collaboration between humans and AI as team members—thereby shaping the future of work with AI. The ultimate objective of these teams must be to achieve complementary team performance (CTP), with the team outperforming each individual team member. The IS community is predestined to drive the development of appropriate theories and to lay the foundation for practical applications. We hope that the conceptual foundation developed in this paper will provide fruitful ground for future research, and that the empirical studies illustrate the validity and potential of the human-AI complementarity paradigm.